### v1

# Lost in Overlap: Exploring Watermark Collision in LLMs

Abstract
The proliferation of large language models (LLMs) in generating content raises concerns about text copyright. Watermarking methods, particularly logit-based approaches, embed imperceptible identifiers into text to address these challenges. However, the widespread use of watermarking across diverse LLMs has led to an inevitable issue known as watermark collision during common tasks like question answering and paraphrasing. This study focuses on dual watermark collisions, where two watermarks are present simultaneously in the same text. The research demonstrates that watermark collision poses a threat to detection performance for detectors of both upstream and downstream watermark algorithms.
Lost in Overlap: Exploring Watermark Collision in LLMs
Yiyang Luo* Nanyang Technological University lawrenceluoyy@outlook.com                        Ke Lin*,† Tsinghua University leonard.keilin@gmail.com                        Chao Gu University of Science and Technology of China guch8017@mail.ustc.edu.cn
footnotetext: *Equal Contribution.footnotetext: Corresponding Author.
1 Introduction
As the text quality produced by large language models (LLM) advances, it tackles numerous practical challenges while raising many new issues. In particular, LLMs’ widespread generation of text on the Internet may increase rumors and raise concerns about text copyright Megías et al. (2021); Tang et al. (2023). Consequently, the identification and classification of machine-generated text have become critically significant. Watermarking methods can address this issue, making them increasingly prominent in current discussions and gathering growing attention.
Text watermarking involves embedding distinctive, imperceptible identifiers (watermarks) into written content. Nowadays, most methods are logit-based Kirchenbauer et al. (2023a); Liu et al. (2023); Zhao et al. (2023); Kuditipudi et al. (2023); Hu et al. (2023): they manipulate the output logits of LLMs during the generation process to achieve the goal of embedding watermarks. These watermarks are sophisticatedly designed to be robust yet discreet, ensuring content integrity and ownership preservation without compromising readability or meaning.
As LLMs flourish, the demand for watermarking rapidly escalates, leading to the application of various watermarking techniques across different LLMs. Consequently, particularly during common tasks such as question answering (QA) and paraphrasing, watermark collision becomes inevitable.
Figure 1: Illustration of watermark collisions.
The term watermark collision can be defined as instances of text with multiple watermarks occurring simultaneously (Fig. 1). While many methods Liu et al. (2023); Zhao et al. (2023); Kuditipudi et al. (2023) claim resilience against paraphrase attacks, none have been specifically tested for watermark collisions. As part of this study, we examine dual watermark collisions in which only two watermarks are involved in the same textual sample.
Figure 2: The method pipeline. $T_{raw}$, $T_{W}$, and $T_{P}$ denote the original input text data, text data with a single watermark, and text with a dual watermark, respectively. For a thorough comparison, we also introduce $T_{P}^{\prime}$, containing only a single watermark. $D_{W}$ and $D_{P}$ represent the watermark detectors for the watermarker $W$ and paraphraser $P$, respectively.
Our Contributions.
The goal of this paper is to discuss watermark collisions and their impact on LLMs. Our contributions are as follows:
  • We propose a novel variant of paraphrase attacks that could easily remove existing watermarks from the text.
  • We find that the strength of the overlapping watermark impacts detection performance. Upstream and downstream watermarks are generally competing for detection accuracy, with one being stronger and the other weaker.
  • We discuss the vulnerability of watermarking techniques caused by watermark collisions.
2 Related Work
Watermarks in the LLM era can be categorized into two categories Liu et al. (2024): post-processing and generation-based techniques. While post-processing techniques Abdelnabi and Fritz (2020); Yang et al. (2022) modify the text and preserve its semantics, generation-based methods incorporate watermarks into the text generation process to achieve better results. Kirchenbauer et al. (2023a) proposed a logit-based greenlist mechanism based on prior token hashes. Liu et al. (2023) introduced a watermark model to generate semantic-preserving logits during text generation. Zhao et al. (2023) simplified the Kirchenbauer et al.’s scheme by using a fixed Green-Red split and achieved greater robustness. Christ et al. (2023) and Kuditipudi et al. (2023) aims to design watermark techniques that are more robust and secure.
Even though these methods have been designed to be more robust against such paraphrase attacks Kirchenbauer et al. (2023b), these paraphrase attacks often use unwatermarked LLMs, e.g. DIPPER Krishna et al. (2023) and GPT-3.5 Brown et al. (2020). Therefore, it is unclear what impact one watermark technique has over another, which is the motivation and objective of this study.
3 Method
We design a pipeline consisting of a watermarker and a paraphraser to simulate the collisions between different watermarks:
  • Watermarker $W$ creates watermarked texts $T_{W}$ by utilizing an LM to generate content based on specific corpus as context.
  • Paraphraser $P$ paraphrases the watermarked texts $T_{W}$ with different watermarks, i.e. generated by different methods or keys.
As shown in Fig. 2, we first generate text data $T_{W}$ using Watermarker $W$. Then the text data will be paraphrased by paraphraser $P$ with different watermarks to generate paraphrased text data $T_{P}$. Paraphrased texts $T_{P}$ are intended to contain dual watermarks simultaneously, and the impact of watermark collisions is then investigated.
Furthermore, we also generate unwatermarked texts $T_{W}^{\prime}$ and $T_{P}^{\prime}$ as baselines using watermarkers and paraphrasers. Section 4.1 and Appendix A provide details regarding pipeline setup.
$W$ $P$ $\text{KGW}_{weak}$ $\text{SIR}_{weak}$ $\text{PRW}_{weak}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{weak}$ 52.80 19.80 3.40 90.09 41.10 48.00
$\text{SIR}_{weak}$ 55.74 25.10 / / 41.05 19.80
$\text{PRW}_{weak}$ 37.00 28.20 22.30 79.56 26.60 41.20
(a) weak $W$, weak $P$
$W$ $P$ $\text{KGW}_{strong}$ $\text{SIR}_{strong}$ $\text{PRW}_{strong}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{weak}$ 4.10 97.40 0.20 92.77 6.20 99.90
$\text{SIR}_{weak}$ 12.45 96.70 / / 13.22 97.50
$\text{PRW}_{weak}$ 14.60 96.70 29.40 91.91 9.30 99.90
(b) weak $W$, strong $P$
$W$ $P$ $\text{KGW}_{weak}$ $\text{SIR}_{weak}$ $\text{PRW}_{weak}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{strong}$ 68.20 21.50 9.00 79.26 56.00 44.40
$\text{SIR}_{strong}$ 61.55 21.00 / / 45.58 14.00
$\text{PRW}_{strong}$ 67.90 18.30 32.00 65.31 44.20 41.70
(c) strong $W$, weak $P$
$W$ $P$ $\text{KGW}_{strong}$ $\text{SIR}_{strong}$ $\text{PRW}_{strong}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{strong}$ 7.00 94.10 0.80 90.52 14.80 99.70
$\text{SIR}_{strong}$ 10.98 97.40 / / 12.58 97.10
$\text{PRW}_{strong}$ 14.60 96.70 37.10 79.00 11.20 99.30
(d) strong $W$, strong $P$
Table 1: TPR of the paraphrased text $T_{P}$ with dual watermarks when $\text{FPR}=1\%$. $W$ and $P$ represent the watermarker and paraphraser, respectively. $D_{W}$ and $D_{P}$ represent the detector of watermarker and paraphraser.
$W$ $T\&D$ ($T_{W}$, $D_{W}$) ($T_{P}^{\prime}$, $D_{W}$) ($T_{P}$, $D_{W}$)
$\text{KGW}_{weak}$ 99.90 71.65 52.80
$\text{KGW}_{strong}$ 100.00 81.95 68.20
$\text{SIR}_{weak}$ 87.90 59.05 55.74
$\text{SIR}_{strong}$ 94.30 67.60 61.55
$\text{PRW}_{weak}$ 95.40 49.25 37.00
$\text{PRW}_{strong}$ 99.70 75.05 67.90
Table 2: TPR of detecting watermark using $D_{W}$ with paraphrase attacks when $\text{FPR}=1\%$. ($T_{W}$, $D_{W}$) represents the TPR before paraphrasing attacks. Meanwhile, ($T_{P}^{\prime}$, $D_{W}$) denotes the TPR after paraphrasing attacks without watermarks, and ($T_{P}$, $D_{W}$) denotes the TPR after paraphrasing attacks with a watermark.
4 Experiments
4.1 Experiment Setup
Settings.
We utilize the C4 dataset Raffel et al. (2020) as the context for text generation with a maximum of 128 tokens. In the capacity of the watermarker, we employ both LLaMA-2-13B Touvron et al. (2023) and OPT-1.3B Zhang et al. (2022), while for paraphrasing, we exclusively utilize LLaMA-2-13B. Appendix A shows more details of the experiments.
Watermarks.
Our experiments were conducted using several previous watermark methods as baselines. The watermark strength of each method is tuned separately for weak and strong watermarks. Methods used include: (a) KGWfrom Kirchenbauer et al. (2023a), $d=2,5$ for weak and strong settings; (b) SIRfrom Liu et al. (2023), $d=2,5$ for weak and strong settings; (c) PRWfrom Zhao et al. (2023), $s=2,5$ for weak and strong settings.
Some methods are not selected for specific reasons. UBW from Hu et al. (2023) asserting that their approach lacks resilience against paraphrase attacks, thereby rendering it unsuitable for this experiment. RDW from Kuditipudi et al. (2023) may be ineffective when subjected to a key different from their recommended configuration, thereby failing to meet the experimental requirements that necessitate distinct key settings.
4.2 Evaluation Metrics
Given the rapid development of the Watermarking field, there is currently no consensus on metrics, with different methods employing varied evaluation criteria. Moreover, many approaches utilize a detection threshold to categorize text as watermarked. Implementing all metrics for each method is impractical and biased. Following the approach of Zhao et al. (2023), we opt for a fair evaluation by avoiding the influence of threshold settings. Our final detection metrics include false positive rates (FPR) and true positive rates (TPR). We specifically set FPR values at 1%, 5%, and 10%, adjusting the detector’s thresholds accordingly. This ensures a consistent and unbiased assessment of watermarking methods. We only present $\text{FPR}=1\%$ here and refer to Appendix B for more results.
4.3 Experimental Results
Paraphrase attacks with watermarks tend to be stronger than those without.
In Table 2, we present the detection accuracy for various baseline watermark algorithms with and without the occurrence of watermark collision ($T_{P}$ and $T_{P}^{\prime}$, respectively). A noteworthy decline in detection accuracy is observed when watermarks are introduced in the context of paraphrase attacks.
Different watermarks overlapping impair detection accuracy, with one being stronger and the other weaker.
It is discernible that the watermark collision adversely impacts the integrity of both the watermarker and the paraphraser, leading to observable damage in their respective watermarks. According to the settings in Table 1, there is a strong competition between overlapping watermarks. As the watermarker attempts to maintain its detection accuracy, the paraphraser’s detection accuracy decreases. SIR-SIR is not listed since the SIR does not allow the user to choose a key to create a different distribution, and therefore the watermark collision will not occur.
It should also be noted that different watermarking methods behave differently during competition. KGW appears to be less competent than the other two methods. SIR, however, shows significant collisions even in weak paraphraser settings (Column SIR of Tab. 0(a) & 0(c)), while PRW exhibits extreme collisions in strong paraphraser settings (Column PRW of Tab. 0(b) & 0(d)).
Upstream watermarks are prone to be erased by downstream watermarks.
The degradation of the original watermark becomes more pronounced with the increasing strength of the paraphrase watermark, as indicated in Tables 0(b) and 0(d). Regardless of the watermarker’s strength, there is a substantial decrease in the detection accuracy of the paraphrased text $T_{P}$ (observed in the $D_{W}$ columns of Table 1). Upstream watermarks experience heightened vulnerability as the downstream paraphraser strengthens, rendering them nearly indistinguishable from unwatermarked text.
If downstream watermarker has a stronger base model than the upstream watermarker, upstream watermarks can be easily erased.
This observation is supported by the use of LLaMA-2-13B and OPT-1.3B as the base models for the watermarker while maintaining LLaMA-2-13B as the base model for the paraphraser, as illustrated in Fig 3. The findings suggest that watermark collision is inevitable across model sizes. Still, the impact of paraphrase attacks becomes more pronounced when the paraphraser’s base model significantly outperforms the watermarker’s.
Figure 3: TPR for watermark detection with various base models for the watermarker, while employing LLaMA-2-13B as the base model for the paraphraser. FPR is set to 1% and both the watermarkers and paraphrasers are configured with weak settings.
5 Discussion
Interference with downstream tasks caused by watermark collisions.
Since LLM watermarks may be used in a variety of downstream applications, particularly in QA scenarios, collisions between watermarks may cause greater damage to downstream watermarks. Texts with watermarks are more likely to be used as user input or dialogue context. As a result, LLM service providers may have difficulty watermarking their output. This situation may occur more frequently in the future and harm the development of LLM watermarks.
Malicious paraphrase attacks based on watermark collisions.
Previous works Kirchenbauer et al. (2023b, a); Kuditipudi et al. (2023) have introduced several attacks, such as copy-paste attacks and paraphrase attacks, but most have shown their robustness and security against at least some of these attacks. Our study, however, provides a feasible method of constructing effective paraphrase attacks using watermark collisions. The use of paraphrases with strong watermarks could easily wipe out existing watermarks, resulting in greater vulnerability to watermarking techniques. This also presents a significant challenge for watermarking techniques development in the future.
Detection of existing watermarks using collisions between watermarks of different strengths.
In Table 2, we demonstrate that weak watermarks can still be easily applied to unwatermarked text. It is, however, much more difficult to apply a weak watermark to text that has already been watermarked (Tab. 0(a) & 0(c)). This provides a simple probabilistic method of detecting watermarks without the need to know their details. When it is difficult to add a weak watermark to a paraphrased sentence, it is more likely to have an existing watermark in the original sentence. Thus, watermark collisions could pose a non-negligible threat to the security of watermarking systems.
6 Conclusion
This paper describes the phenomenon of watermark collisions, whereby overlapping watermarks in the same text can reduce the accuracy of both upstream and downstream watermark detection. We conduct our experiments by combining various watermarkers and paraphrasers. In our study, we found that watermark collisions result in overlapping watermarks competing with each other, with one watermark being stronger than the other. We further highlight that watermark collisions may jeopardize the validity and security of watermarks. While we do not intend to resolve this problem of watermark collisions, we hope our work may raise awareness of potential threats to LLM watermarking.
Limitations
Our approach indeed demonstrates the potential collision between existing watermark techniques. Nevertheless, we conduct our experiments only on paraphrasing tasks to simulate watermark collisions. Broader tasks, such as question answering, have not yet been tested.
Furthermore, we select the LLaMA-2-13B as the paraphraser, whereas some other models (e.g. DIPPER) are specially designed for paraphrasing. We believe that the LLaMA has shown its generalization ability in a variety of scenarios. This leads us to believe that it is also capable of paraphrasing.
Also, the maximum number of generated tokens is set to 128. Our research suggests that the observation of watermark collisions would also be evident in a longer context since all logit-based watermarks are susceptible to destructive logit shifts like this. Further experiments are still required.
Acknowledgement
Appendix A Pipeline Setup
Datasets.
To generate the watermarked text $T_{W}$, the C4 dataset is used as the context. A total of 1000 watermarked textual samples are generated from the selected context. To ensure that only text with an apparent watermark is selected for $T_{W}$, a z-score threshold is set during the generation. For KGW and PRW, the z-threshold is set to $4.0$. For SIR, the z-threshold is set to $0.0$.
Hyperparameters.
We specifically designate $2024$ as the key for the watermarker (if applicable) and $2023$ as the key for the paraphraser (if applicable) to vary the key for watermark collision within the same watermark algorithm (e.g. KGW-KGW). The specific hyperparameters of each watermarking method are as follows:
  • KGW: For weak settings, we set the green list size $\gamma=0.25$, hardness parameter $\delta=2.0$ and the seeding scheme is selfhash. For strong settings, we set the green list size $\gamma=0.25$, hardness parameter $\delta=5.0$ and the seeding scheme is selfhash.
  • SIR: We employ the context mode. For weak settings, We set $chunk\_length=10$, $\gamma=0.5$, watermark strenght $\delta=2.0$. For strong settings, we set $chunk\_length=10$, $\gamma=0.5$, watermark strength $\delta=5.0$.
  • PRW: For weak settings, we set the green list size $\gamma=0.25$, watermark strength $\delta=2.0$. For strong settings, we set the green list size $\gamma=0.25$, watermark strength $\delta=5.0$.
Prompts.
We formulate a prompt tailored for LLaMA-2-13B, enabling it to proficiently paraphrase the given content. The prompt template is shown in Fig. 4.
<<SYS>>
Assume you are a helpful assistant. Your job is to paraphase the given text.
<</SYS>>
[INST] {INPUT_TEXT}[/INST]
You’re welcome! Here’s a paraphrased version of the original message:
Figure 4: The paraphrase template for paraphraser.
Appendix B Experimental Results
Table 3 shows the TPR results of detection when utilizing LLaMA-2-13B as the base model of watermarker. Table 4 shows the TPR results of detection when utilizing OPT-1.3B as the base model of watermarker.
Watermarker $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
Paraphraser $\text{KGW}_{weak}$ $\text{KGW}_{strong}$ $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 52.80 19.80 4.10 97.40 68.20 21.50 7.00 94.10
5% 66.60 34.20 14.30 99.60 72.80 32.70 16.20 97.90
10% 71.00 41.50 22.10 99.60 75.00 41.70 26.10 99.40
Watermarker $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
Paraphraser $\text{SIR}_{weak}$ $\text{SIR}_{strong}$ $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 3.40 90.09 0.20 92.77 9.00 79.26 0.80 90.52
5% 5.90 92.12 2.00 94.66 14.10 89.41 2.50 93.76
10% 7.80 93.81 3.80 95.88 17.30 92.14 4.70 95.88
Watermarker $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
Paraphraser $\text{PRW}_{weak}$ $\text{PRW}_{strong}$ $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 41.10 48.00 6.20 99.90 56.00 44.40 14.80 99.70
5% 54.40 65.70 16.70 99.90 64.70 64.30 25.40 100.00
10% 60.20 71.70 23.70 99.90 68.50 70.80 33.00 100.00
Watermarker $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
Paraphraser $\text{KGW}_{weak}$ $\text{KGW}_{strong}$ $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 55.74 25.10 12.45 96.70 61.55 21.00 10.98 97.40
5% 65.97 34.80 25.86 99.20 72.95 30.90 27.56 98.80
10% 69.73 40.30 31.22 99.50 75.44 38.70 34.23 99.40
Watermarker $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
Paraphraser $\text{PRW}_{weak}$ $\text{PRW}_{strong}$ $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 41.05 19.80 13.22 97.50 45.58 14.00 12.58 97.10
5% 52.67 56.60 22.43 100.00 59.98 45.90 25.48 99.70
10% 56.17 67.30 27.63 100.00 63.17 59.00 30.49 99.90
Watermarker $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
Paraphraser $\text{KGW}_{weak}$ $\text{KGW}_{strong}$ $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 37.00 28.20 14.60 96.70 67.90 18.30 14.60 96.70
5% 56.40 38.60 22.30 99.20 73.50 29.70 22.30 99.20
10% 65.80 45.70 30.20 99.40 76.50 38.70 30.20 99.40
Watermarker $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
Paraphraser $\text{SIR}_{weak}$ $\text{SIR}_{strong}$ $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 22.30 79.56 29.40 91.91 32.00 65.31 37.10 79.00
5% 31.80 91.00 37.70 94.43 42.20 87.78 47.10 90.89
10% 39.20 92.56 42.10 95.63 47.90 91.19 51.20 93.22
Watermarker $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
Paraphraser $\text{PRW}_{weak}$ $\text{PRW}_{strong}$ $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 26.60 41.20 9.30 99.90 44.20 41.70 11.20 99.30
5% 37.80 67.20 14.30 100.00 50.60 58.50 18.60 99.90
10% 45.30 72.00 18.60 100.00 55.50 65.20 24.90 99.90
Table 3: TPR of the paraphrased text $T_{P}$ with dual watermarks when utilizing LLaMA-2-13B as the base model of watermarker. FPR is set to $1\%$, $2\%$ $5\%$, respectively.
Watermarker $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
Paraphraser $\text{KGW}_{weak}$ $\text{KGW}_{strong}$ $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 1.40 23.40 4.30 98.00 1.50 26.30 6.90 97.90
5% 10.90 34.30 17.60 99.40 7.30 36.50 19.90 99.20
10% 17.20 44.10 25.60 99.60 15.30 44.20 28.30 99.40
Watermarker $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
Paraphraser $\text{SIR}_{weak}$ $\text{SIR}_{strong}$ $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 0.80 92.81 0.90 85.32 0.90 72.27 1.20 87.80
5% 3.60 94.72 3.80 91.24 3.80 87.62 4.10 93.63
10% 6.70 96.18 6.70 93.65 7.40 91.60 8.10 95.49
Watermarker $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
Paraphraser $\text{PRW}_{weak}$ $\text{PRW}_{strong}$ $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 5.00 36.80 6.90 99.80 5.80 14.70 8.80 97.40
5% 21.10 56.10 22.30 100.00 17.80 52.30 22.00 99.80
10% 34.30 66.20 31.60 100.00 29.00 63.00 31.50 100.00
Watermarker $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
Paraphraser $\text{KGW}_{weak}$ $\text{KGW}_{strong}$ $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 1.25 19.30 5.17 96.40 2.61 17.00 6.52 93.70
5% 4.05 32.30 15.82 98.30 6.58 26.90 16.04 98.80
10% 6.76 40.50 20.34 98.70 10.44 35.20 20.32 99.30
Watermarker $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
Paraphraser $\text{PRW}_{weak}$ $\text{PRW}_{strong}$ $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 2.52 27.20 4.26 97.70 3.76 13.70 8.25 97.80
5% 5.45 52.20 10.93 99.70 8.57 46.00 13.14 100.00
10% 8.28 59.70 13.77 99.80 11.60 55.60 15.53 100.00
Watermarker $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
Paraphraser $\text{KGW}_{weak}$ $\text{KGW}_{strong}$ $\text{KGW}_{weak}$ $\text{KGW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 0.10 23.50 5.70 98.60 1.30 17.50 7.80 96.80
5% 2.90 35.60 17.30 99.40 6.30 29.70 21.80 98.60
10% 13.60 44.20 34.00 99.50 11.70 42.10 29.60 98.70
Watermarker $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
Paraphraser $\text{SIR}_{weak}$ $\text{SIR}_{strong}$ $\text{SIR}_{weak}$ $\text{SIR}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 10.60 87.23 24.90 88.96 14.00 78.84 24.30 88.62
5% 24.30 91.38 38.60 93.93 30.30 89.81 38.50 93.53
10% 38.90 92.50 48.60 94.92 36.60 92.39 43.60 94.53
Watermarker $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
Paraphraser $\text{PRW}_{weak}$ $\text{PRW}_{strong}$ $\text{PRW}_{weak}$ $\text{PRW}_{strong}$
FPR $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
1% 0.80 36.10 4.70 99.60 1.00 11.60 6.10 93.30
5% 3.70 57.80 10.70 100.00 6.50 52.10 13.80 100.00
10% 10.30 66.00 19.90 100.00 11.60 63.20 18.40 100.00
Table 4: TPR of the paraphrased text $T_{P}$ with dual watermarks when utilizing OPT-1.3B as the base model of watermarker. FPR is set to $1\%$, $2\%$ $5\%$, respectively.
Appendix C Scientific Artifacts
The licenses for all the watermarking methods are listed below: KGW (Apache 2.0 Licence), SIR (MIT Licence), PRW (MIT Licence). The licenses for models are listed below: LLaMA-2-13B (LLAMA 2 Community License), OPT-1.3B (OPT LICENSE).